# Neural Interpretable Reasoning

### v1

@@ 1 A Turing test for interpretability
Donal Duck attempts to start his car, model 313, but the vehicle fails to start. After inspecting the situation, he finds that the fuel level is too low. Once he refuels, the car starts without issue. In this instance, Donal clearly understands the problem and its straightforward solution. The following day, the car again fails to start despite a full fuel tank. Uncertain of the cause, Donal consults a mechanic. The mechanic, with his expertise in engines, diagnoses an oil leak as the root of the problem. After repairing the leak, the car operates normally. Here, while Donal could not diagnose the issue on his own, his recourse to expert knowledge ultimately resolved the problem.
These examples illustrate that understanding a system is often subjective and dependent on the user’s background (Miller, 2019). However, they also suggest a practical criterion to check whether a system is interpretable for a user which can informally present as follows:
@@
The purpose of this work is threefold:
@@
  • Break combinatorial complexity in verifying interpretability: We show how the combinatorial complexity in verifying inference equivariance can be mitigated considering interpretability as a Markovian property and using technques such as neural re-parametrization and mixture models (Sec. 3).
@@ 2 Interpretability & equivariance
Our work is motivated by the idea that a system is interpretable if its internal processes can be reliably translated into outcomes that users can predict. In this section, we formalize this notion as interpretability equivariance, establishing that performing inference using the system’s mechanisms should commute with the process of inference performed using the user’s mechanisms, as illustrated in the following example.
@@
Donald Duck wants to sleep but is uncomfortably cold. To achieve a comfortable sleep, he needs to warm up his environment to an appropriate temperature. The heating system is controlled by a thermostat, yet Donald has misplaced the user manual. The thermostat provides only two pieces of information: a wheel with 8 positions (currently set to $1$) and a numeric display ranging from 0 to 10 (currently showing $3$).
@@
Donal Duck path: $\quad\texttt{wheel}=4\to\texttt{heat}=\text{medium}\to\texttt{comfort}=\text{yes}.$
The example illustrates that while the thermostat’s variables differ in semantics from Donald Duck’s internal concepts, they are nonetheless aligned closely enough for him to establish a straightforward mapping between the two. For instance, a wheel position within the range $[3,4]$ might be interpreted as medium heat, and a display reading of $2$ may be associated with a state of comfort. Furthermore, Donald’s reasoning process demonstrates that he can deduce the system’s state via two equivalent routes–either by consulting the display or by directly sensing the heat output–with both methods leading to the same conclusion. In the next paragraph we introduce the notation we will use to formalize this equivariance.
@@ 2.1 Interpretability as inference equivariance
Let $V$ denote a set of random variables representing different aspects of a system (for example, heating levels, wheel position, etc.) with a joint probability distribution
$\mathbb{P}(V)=\mathbb{P}(V_{1},V_{2},\ldots,V_{n}).$
In many scenarios, we are interested in the conditional probability of a particular query variable $V_{q}$ (such as the comfort level) given some evidence $V_{e}$ (like the wheel position), denoted by $\mathbb{P}(V_{q}\mid V_{e})$. To formalize the distinction between the internal, machine-oriented description of the system and its human-interpretable counterpart, we index machine-related variables with the superscript $m$ (so that $V^{(m)}$ represents the machine’s variables) and human-related variables with $h$. A translation function
$\tau:V^{(m)}\to V^{(h)}$
is introduced to map the machine variables into the human reference system. Consequently, for any distribution $\mathbb{P}(V^{(m)})$ over the machine variables, the corresponding distribution in the human space is given by the push-forward measure:
$\mathbb{P}\bigl{(}\tau(V^{(m)})\bigr{)}=\tau\bigl{(}\mathbb{P}(V^{(m)})\bigr{)}.$
This expression means that applying the translation function to the machine’s probability distribution produces a distribution over the human-interpretable variables.
@@
The principle of inference equivariance, illustrated in the Donald Duck’s example, asserts that the process of translating a machine’s probability distribution into the human reference system and then querying it should yield the same result as first performing the query within the machine’s domain and then translating the result. Formally, this is expressed as
$\mathbb{P}\bigl{(}\tau(V^{(m)}_{q})\mid\tau(V^{(m)}_{e})\bigr{)}=\tau\Bigl{(}\mathbb{P}\bigl{(}V^{(m)}_{q}\mid V^{(m)}_{e}\bigr{)}\Bigr{)}.$
This equality encapsulates the idea that whether one chooses to “translate then query” or “query then translate”, the resulting inference remains consistent, as already observed for causal structures (Geiger et al., 2024; Marconato et al., 2023). In the context of the Donal Duck example, this principle becomes particularly clear. Donal Duck faces a thermostat whose internal variables—such as the wheel setting and display reading—are not immediately aligned with his intuitive notions of heat and comfort [^1] Notice that in contrast with equivariances in causal abstractions (Geiger et al., 2024) where the inference structure is assumed to be aligned with the true data generating mechanisms.. By establishing a mapping between the machine’s outputs and his own reference system, he is able to reliably predict his comfort level.
For instance, Donal might first translate the thermostat’s raw signal (the display reading) into his internal concept of temperature and then infer his comfort state based on that interpretation. Alternatively, he might directly observe the mechanical behavior (the wheel position) to predict the corresponding change in room temperature, and only afterward translate that information into his subjective experience of warmth. The fact that both routes lead him to the same conclusion—whether he “translates then queries” or “queries then translates”—demonstrates the principle of inference equivariance.
This consistency is critical: it ensures that the mapping between machine variables and human concepts is robust, thereby making the system interpretable. In essence, the equality translate, then query = query, then translate guarantees that the user’s understanding and predictions of the system’s behavior remain coherent, regardless of the order in which translation and inference occur.
@@
While the concept of equivariance provides a robust framework for linking machine and human perspectives, its practical implementation is fraught with challenges. As the number of variables increases, verifying and maintaining equivariance becomes exponentially more complex. To illustrate, consider a simple scenario where every variable in the system is Boolean. In this simple case, a complete interpretation of the system would require to verify the equivariance for all possible states of the system. This corresponds to extract the full conditional probability table, which contains $2^{n}$ entries for $n$ variables. Even for a modest $n$, the number of combinations quickly becomes computationally intractable. For this reason, in practical applications it becomes essential to guarantee inference equivariance indirectly or approximately, while maintaining computational efficiency. In practice, this may involve constraining the inference space to a subset of critical variables, leveraging problem-specific structures to reduce complexity, or employing surrogate models that approximate the full system behavior with significantly lower computational cost.
@@ 2.2 Properties of interpretability through the lenses of inference equivariance
Inference equivariance can be asymmetric: In the thermostat example, Donal Duck uses the available signals—such as the wheel position and the display reading—to form an understanding of the system’s behavior. Importantly, for him to use the thermostat effectively, it is not necessary to have a complete, invertible mapping from his internal concepts (e.g., comfort level) back to the machine’s variables. This one-way, or asymmetric, mapping suffices because Donal only needs to translate machine outputs into human-understandable signals. The absence of a reverse transformation does not impede his ability to predict the system’s response, illustrating that the forward mapping (machine to human) is all that is required for interpretability (but the opposite might be needed for supervised learning).
Explanations are a form of selection: An explanation of a system’s behavior can be seen as a process of selection, where conditioning on observed evidence picks out a specific subset from the system’s complete conditional probability table. In the Donald Duck example, when Donald observes a particular display reading or wheel position, he effectively selects a corresponding segment of the conditional probability table that relates these inputs to his comfort state. This selection—formally represented as $\mathbb{P}(V\mid V^{\prime}=v^{\prime})$—encapsulates the explanation by narrowing down the myriad potential outcomes to the ones relevant to his observation.
@@
Post-hoc methods complicate rather than simplify interpretability: When applying post-hoc interpretability techniques, such as using surrogate models to explain the original system, an additional layer of equivariance is required. Suppose Donald employs a surrogate model to better understand his thermostat. In that case, there must be a consistent mapping between the machine variables of the original system and those of the surrogate model and another mapping from the surrogate model to Donald Duck. Formally, if we denote the translation and inference functions for the surrogate as $\tau^{\prime}$ and $\eta^{\prime}$ respectively, both the original and surrogate systems must satisfy equivariance conditions:
$\eta^{\prime}\Bigl{(}\tau^{\prime}\bigl{(}\mathbb{P}(V^{(m)})\bigr{)}\Bigr{)}=\tau^{\prime}\Bigl{(}\eta^{\prime}\bigl{(}\mathbb{P}(V^{(m)})\bigr{)}\Bigr{)}$
and
$\eta\Bigl{(}\tau\bigl{(}\mathbb{P}(V^{(m^{\prime})})\bigr{)}\Bigr{)}=\tau\Bigl{(}\eta\bigl{(}\mathbb{P}(V^{(m^{\prime})})\bigr{)}\Bigr{)}.$
This requirement ensures that the explanations generated by the surrogate model faithfully reflect the behavior of the original system, thus preserving interpretability even when using post-hoc methods. Ultimately, the need to establish these additional mappings significantly complicates the interpretability process as two inference equivariances must be satisfied instead of one.
@@ 2.3 Semantic and functional equivariances
Previous works (Geiger et al., 2024; Marconato et al., 2023) focused primarily on semantic equivariance, emphasizing that equivariance should hold for random variables $V$. However, less attention has been paid to the functions that describe mappings between random variables; for a user to truly understand the underlying mechanisms, the structure of the function and its parameters must also satisfy equivariance, as illustrated in the following example.
@@
$\mathbb{P}(V_{2}\mid V_{1};\mu,\sigma)\coloneqq\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\!\left(-\frac{(v-\mu)^{2}}{2\sigma^{2}}\right).$
@@ 3.1 Interpretability is a Markovian property
In the earlier thermostat example, Donald Duck successfully built an intuitive understanding of how the thermostat worked, despite having no specialized knowledge of electronics or physics. This observation illustrates how interpretability is a Markovian property: a user can interpret a system at a given level of abstraction without needing to reference lower-level details. In this context, interpretability is achieved locally—each step of the inference process can be understood in isolation from others. Formally, we can formalize this Markovian property of interpretability by writing:
$\mathbb{P}\models(V_{i}\bot V\notin\text{pa}(V))\mid\text{pa}(V),$
meaning that any variable $V_{i}$ is conditionally independent of all non-parent variables, given its parent variables. This property allows a user to interpret a single step of the inference process—the one concerning the variable $V_{i}$—without needing to backtrack through the entire chain of reasoning.
This Markovian property of interpretability attenuates scalability issues, as it permits the analysis of individual steps without the burden of interpreting the entire system at once. This layered approach is reflected in models such as Concept Bottleneck Models (Koh et al., 2020), Neural Additive Models (Agarwal et al., 2021), Decision Trees (Breiman et al., 1984), or Prototypical Networks (Chen et al., 2019), where each layer or component is designed to be interpretable on its own. In the Donald Duck example, his ability to understand the thermostat’s behavior without delving into the intricate details of its engineering underscores the practical benefits of this Markovian property.
@@
This example highlights a fundamental scalability issue: while a machine can, in principle, process and manage a large number of independent variables, human users typically can only handle around 7 $\pm$ 2 variables at any one time (Aristotle, (n.d.; Miller, 1956). It clear that even under the assumption that variables operate independently (which is quite common in XAI), the number of interactions required to understand the system grows linearly with the number of variables. For humans, who are limited to processing a constant (i.e., 7 ± 2) number of variables simultaneously, this poses a significant obstacle to interpretability. The key question then becomes: how can we design a system that presents only a constant number of variables to a human, without sacrificing the system’s overall expressivity? A promising approach to manage this challenge is re-parametrization, where a system is transformed into an equivalent form that preserves its expressivity while reducing the number of variables that a human must directly consider.
@@
One effective strategy is to decompose a complex system into a mixture of simpler subsystems, each of which is easy to understand (McLachlan & Basford, 1988). For instance, imagine a thermostat with 365 knobs (so, even more than the original 100 knobs!), but with the twist that only one knob is active per day, and an indicator light signals which knob is relevant at that time. This design ensures that, at any given moment, Donald needs to focus on only one knob rather than hundreds. Such re-parametrization retains the full expressive power of the original system while offering local representations that are much more interpretable. Techniques like LIME (Ribeiro et al., 2016), ProtopNets (Chen et al., 2019), and Concept Memory Reasoning (Debot et al., 2024) embody this approach by generating simple, locally faithful explanations whose composition may form arbitrarily non-linear decision boundaries.
@@
In many classification problems, re-parametrization involves two key components: mapping raw variables to higher-level concepts (semantic re-parametrization) and decomposing complex function parameters into simpler mixtures (functional re-parametrization). In this framework, the original data variables are transformed into a set of human-interpretable concepts, ensuring semantic transparency as in Concept Bottleneck Models (Koh et al., 2020). Simultaneously, the function that governs the model’s behavior is restructured into a mixture of simple functions, which preserves the model’s expressivity while making it easier to understand as in Neural Additive Models (Agarwal et al., 2021).

### v2

@@ 1 A Turing test for interpretability
Donald Duck attempts to start his car, model 313, but the vehicle fails to start. After inspecting the situation, he finds that the fuel level is too low. Once he refuels, the car starts without issue. In this instance, Donald clearly understands the problem and its straightforward solution. The following day, the car fails to start once more despite having a full fuel tank. Uncertain of the cause, Donald consults a mechanic. Building on her expertise in engines, the mechanic determines that an oil leak is the root of the problem. After repairing the leak, the car operates normally. Here, while Donald could not diagnose the issue on his own, his recourse to expert knowledge ultimately resolved the problem.
These examples illustrate that understanding a system is often subjective and dependent on the user’s background (Miller, 2019). However, they also suggest a practical criterion to check whether a system is interpretable. We can informally describe this criterion as follows:
@@
This work’s purpose can be characterized as threefold:
@@
  • Break combinatorial complexity in verifying interpretability: We show how the combinatorial complexity in verifying inference equivariance can be mitigated considering interpretability as a Markovian property and using techniques such as neural re-parametrization and mixture models (Sec. 3).
@@ 2 Interpretability & equivariance
Our work is motivated by the idea that a system is interpretable if its internal processes can be reliably translated into outcomes that users can predict. In this section, we formalize this notion as interpretability equivariance, establishing that performing inference using the system’s mechanisms should commute with the process of inference performed using the user’s mechanisms. We begin by motivating and illustrating this definition via an example:
@@
Donald Duck wants to sleep but is uncomfortably cold. To achieve a comfortable sleep, he needs to warm up his environment to an appropriate temperature. A thermostat, whose user’s manual Donald misplaced, controls the heating system. The thermostat provides only two pieces of information: a wheel with eight positions (currently set to $1$) and a numeric display ranging from 0 to 10 (currently showing $3$).
@@
Donald Duck path: $\quad\texttt{wheel}=4\to\texttt{heat}=\text{medium}\to\texttt{comfort}=\text{yes}.$
The example illustrates that while the thermostat’s variables differ in semantics from Donald Duck’s internal concepts, they are nonetheless aligned closely enough for him to establish a straightforward mapping between the two. For instance, a wheel position within the range $[3,4]$ might be interpreted as medium heat, and a display reading of $2$ may be associated with a state of comfort. Furthermore, Donald’s reasoning demonstrates that he can deduce the system’s state via two equivalent routes–either by consulting the display or by directly sensing the heat output–with both methods leading to the same conclusion. Building on this intuition, we first introduce some useful notation and then use this to formalize our notion of interpretability equivariance.
@@ 2.1 Interpretability as inference equivariance
Let $V$ denote a set of random variables representing different aspects of a system (for example, heating levels, wheel position, etc.). We write the joint probability distribution of these variables as $\mathbb{P}(V)=\mathbb{P}(V_{1},V_{2},\cdots,V_{n}).$ To formalize the distinction between the internal, machine-oriented description of the system and its human-interpretable counterpart, we index machine-related variables with the superscript $m$ (so that $V^{(m)}$ represents the machine’s variables) and human-related variables with $h$. Following Rubenstein et al. (2017), we define a translation function $\tau:V^{(m)}\to V^{(h)}$ as a map between machine variables and the variables within the human’s reference system. Consequently, for any distribution $\mathbb{P}(V^{(m)})$ over the machine variables, the corresponding distribution in the human space is given by the push-forward measure $\mathbb{P}_{\tau(V^{(m)})}=\tau\bigl{(}\mathbb{P}_{V^{(m)}}\bigr{)}$. In particular, for each action on the $i$-th machine variable $a(i)$ (e.g., observing $a(i)\coloneqq\bigl{(}V_{i}^{(m)}=k\bigr{)}$ or intervening on the value of a variable $a(i)\coloneqq do(V_{i}^{(m)})$), we can define the induced distribution $\mathbb{P}_{\tau(V^{(m)})}^{i}=\tau\bigl{(}\mathbb{P}_{V^{(m)}}^{a(i)}\bigr{)}$. To exactly transform the machine system into the human system, we require a surjective mapping $\omega:I_{V^{(m)}}\to I_{V^{(h)}}$ that assigns machine variable indices to human variable indices such that $\mathbb{P}_{\tau(V^{(m)})}^{i}=\mathbb{P}_{V^{(h)}}^{a(\omega(i))}$. Rather than enforcing that $\omega$ be order-preserving as in Rubenstein et al. (2017), our formulation of inference equivariance requires that $\omega$ preserves conditional independence relations (which represents a weaker requirement). Formally, define the neighborhood of a machine variable $V_{i}^{(m)}$ as the minimal set of variables rendering it conditionally independent of the rest, i.e.,
$\mathcal{N}(V_{i}^{(m)})=\min\{\,S^{(m)}\subseteq V^{(m)}\setminus\{V_{i}^{(m)}\}:V_{i}^{(m)}\perp(V^{(m)}\setminus(\{V_{i}^{(m)}\}\cup S^{(m)}))\mid S^{(m)}\,\}.$
We say that $\omega$ preserves conditional independencies if and only if, for every $V_{i}^{(m)}$ and every subset $S^{(m)}\subseteq V^{(m)}\setminus\{V_{i}^{(m)}\}$,
$V_{i}^{(m)}\perp(V^{(m)}\setminus(\{V_{i}^{(m)}\}\cup S^{(m)}))\mid S^{(m)}$
if and only if
$\tau(V_{\omega(i)}^{(m)})\perp(\tau(V^{(m)})\setminus(\{\tau(V_{\omega(i)}^{(m)})\}\cup\tau(S^{(m)})))\mid\tau(S^{(m)}).$
This condition ensures that $\omega$ precisely mirrors the conditional independence structure between the machine and human systems.
@@
The principle of inference equivariance, illustrated in our previous example, asserts that the process of translating a machine’s probability distribution into the human reference system and then querying it should yield the same result as first performing the query within the machine’s domain and then translating the result. Formally, this is expressed as
[…]
This equality encapsulates the idea that whether one chooses to “translate, then query” or to “query, then translate”, the resulting inference remains the same, as already observed for causal structures (Rubenstein et al., 2017; Geiger et al., 2024; Marconato et al., 2023). In the context of the Donald Duck example, this principle becomes particularly clear. Donald Duck faces a thermostat whose internal variables—such as the wheel setting and display reading—are not immediately aligned with his intuitive notions of heat and comfort [^1] Notice that in contrast with equivariances in causal abstractions (Geiger et al., 2024) where the inference structure is assumed to be aligned with the true data generating mechanisms.. By establishing a mapping between the machine’s outputs and his own reference system, he is able to reliably predict his comfort level.
For instance, Donald might first translate the thermostat’s raw signal (the display reading) into his internal concept of temperature and then infer his comfort state based on that interpretation. Alternatively, he might directly observe the mechanical behavior (the wheel position) to predict the corresponding change in room temperature, and only afterwards translate that information into his subjective experience of warmth. The fact that both routes lead him to the same conclusion—whether he “translates, then queries” or “queries, then translates”—demonstrates the principle of inference equivariance.
This consistency is critical: it ensures that the mapping between machine variables and human concepts is robust, thereby making the system interpretable. In essence, the equality “translate, then query” = “query, then translate” guarantees that a user’s understanding and predictions of a system’s behavior remain coherent, regardless of the order in which translation and inference occur.
@@
While the concept of equivariance provides a robust framework for linking machine and human perspectives, its practical implementation is fraught with challenges. As the number of variables increases, verifying and maintaining equivariance becomes exponentially more complex. To illustrate, consider a simple scenario where every variable in the system is Boolean. In this simple case, a complete interpretation of the system would require verifying the equivariance for all possible states of the system. This corresponds to extracting the full conditional probability table, which contains $2^{n}$ entries for $n$ variables. Even for a modest $n$, the number of combinations quickly becomes computationally intractable. For this reason, in practical applications it becomes essential to guarantee inference equivariance indirectly or approximately while maintaining computational efficiency. In practice, this may involve constraining the inference space to a subset of critical variables, leveraging problem-specific structures to reduce complexity, or employing surrogate models that approximate the full system’s behavior with a significantly lower computational cost.
@@ 2.2 Properties of interpretability through the lenses of inference equivariance
Inference equivariance can be asymmetric: In the thermostat example, Donald Duck uses the available signals—such as the wheel position and the display reading—to form an understanding of the system’s behavior. Importantly, for him to use the thermostat effectively, it is unnecessary to have a complete, invertible mapping from his internal concepts (e.g., “comfort level”) back to the machine’s variables. This one-way, asymmetric mapping suffices because Donald only needs to translate machine outputs into human-understandable signals. The absence of a reverse transformation does not impede his ability to predict the system’s response, illustrating that the forward mapping (machine $\rightarrow$ human) is all we require for interpretability (although the opposite mapping might be needed for supervised learning).
Explanations are a form of selection: An explanation of a system’s behavior can be seen as a process of selection, where conditioning on observed evidence picks out a specific subset from the system’s complete conditional probability table. In the Donald Duck example, when Donald observes a particular display reading or wheel position, he effectively selects a corresponding segment of the conditional probability table that relates these inputs to his comfort state. This selection—formally represented with the distribution $\mathbb{P}(V\mid a(V^{\prime}))$—encapsulates the explanation by narrowing down the myriad potential outcomes to the ones relevant to his observation.
@@
Post-hoc methods complicate rather than simplify interpretability: When applying post-hoc interpretability techniques, such as using surrogate models to explain the original system (Hinton, 2015; Zilke et al., 2016) or so-called feature importance methods (Ribeiro et al., 2016; Lundberg & Lee, 2017; Erhan et al., 2009; Sundararajan et al., 2017), an additional layer of equivariance is required. Suppose Donald employs a surrogate model to better understand his thermostat. In that case, there must be a consistent mapping between the machine variables of the original system $V^{(m)}$ and those of the surrogate model $V^{(s)}$ and another mapping from the surrogate model to Donald Duck $V^{(h)}$. Formally, both the original and surrogate systems must satisfy the inference equivariance conditions:
[…]
This requirement ensures that the explanations generated by the surrogate model faithfully reflect the behavior of the original system, thus preserving interpretability even when using post-hoc methods. Ultimately, the need to establish these additional mappings significantly complicates the interpretability process as two equivariance relations must be satisfied instead of one.
@@ 2.3 Semantic and functional equivariances
Previous works (Geiger et al., 2024; Marconato et al., 2023) focused primarily on semantic equivariance, emphasizing that equivariance should hold for random variables $V$. However, less attention has been paid to the functions that describe the mappings between random variables; for a user to truly understand the underlying mechanisms, the structure of the function and its parameters must also satisfy equivariance, as illustrated in the following example.
@@
$\mathbb{P}(V_{2}=v;\mu=V_{1},\sigma)\coloneqq\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\!\left(-\frac{(v-\mu)^{2}}{2\sigma^{2}}\right).$
@@ 3.1 Interpretability is a Markovian property
In the earlier thermostat example, Donald Duck successfully built an intuitive understanding of how the thermostat worked, despite having no specialized knowledge of electronics or physics. This observation illustrates how interpretability is a Markovian property: a user can interpret a system at a given level of abstraction without needing to reference lower-level details. In this context, interpretability is achieved locally—each step of an inference process can be understood in isolation from others. We can formalize this Markovian property of interpretability by writing:
$\forall\;V_{i},V\;\mathbb{P}\models(V_{i}\bot V)\mid\mathcal{N}(V_{i})$ (1)
meaning that, given its 1-hop neighborhood $\mathcal{N}(V_{i})$, any variable $V_{i}$ is conditionally independent of all other variables. This property allows a user to interpret a single step of the inference process—the one concerning the variable $V_{i}$—without needing to backtrack through the entire chain of reasoning.
This Markovian property of interpretability attenuates scalability issues, as it permits the analysis of individual steps without the burden of interpreting the entire system at once. This layered approach is reflected in models such as Self-Explaining Neural Networks (Alvarez Melis & Jaakkola, 2018), Concept Bottleneck Models (Koh et al., 2020), or Prototypical Networks (Chen et al., 2019), where semantically interpretable components (e.g., the concept bottleneck) are designed to be interpretable on their own, regardless of previous layers. In the Donald Duck example, his ability to understand the thermostat’s behavior without the need to understand its engineering shows the practical benefits of this Markovian property.
@@
This example highlights a fundamental scalability issue: while a machine can, in principle, process and manage a large number of independent variables, human users typically can only handle around 7 $\pm$ 2 variables at any one time (Miller, 1956). It clear that even under the assumption that variables operate independently (which is quite common in the field of eXplainable Artificial Intelligence, or XAI), the number of interactions required to understand the system grows linearly with the number of variables. For humans, who are limited to processing a constant number of variables simultaneously (i.e., 7 ± 2), this poses a significant obstacle to interpretability. The key question then becomes: how can we design a system that presents only a constant number of variables to a human, without sacrificing the system’s overall expressivity? A promising approach to manage this challenge is re-parametrization, where a system is transformed into an equivalent form that preserves its expressivity while reducing the number of variables a human must directly consider.
@@
One effective strategy is to decompose a complex system into a mixture of simpler subsystems, each of which is easy to understand (McLachlan & Basford, 1988). For instance, imagine a thermostat with 365 knobs (so, even more than the original 100 knobs!), but with the twist that only one knob is active per day, and an indicator light signals which knob is relevant at that time. This design ensures that, at any given moment, Donald needs to focus on only one knob rather than hundreds. Such re-parametrization retains the full expressive power of the original system while offering local representations that are much more interpretable. Techniques like Self-Explaining Neural Networks (Alvarez Melis & Jaakkola, 2018), ProtopNets (Chen et al., 2019), and Concept Memory Reasoning (Debot et al., 2024) embody this approach by generating simple, locally faithful explanations whose composition may form arbitrarily non-linear decision boundaries.
@@
In many classification problems, re-parametrization involves two key components: mapping raw variables to higher-level concepts (semantic re-parametrization) and decomposing complex function parameters into simpler mixtures (functional re-parametrization). In this framework, the original data variables are transformed into a set of human-interpretable concepts, ensuring semantic transparency as in Concept Bottleneck Models (Koh et al., 2020). Simultaneously, the function that governs the model’s behavior is restructured into a mixture of simple functions, which preserves the model’s expressivity while making it easier to understand as in Self-Explaining Neural Networks (Alvarez Melis & Jaakkola, 2018) and Concept Memory Reasoning (Debot et al., 2024).